OpenStudy (anonymous):

find the equation of the line with the given properties. Express the equation in general form or in slope intercept form. perpendicular to the line 3x+y=-14; contains the point (-6,3)

OpenStudy (anonymous):

so 3x+y = -14. so y = -3x - 14. if u compare this with y = mx+c , M is the slope. so what is the slope of this line now?